# Improving Instruction-Following in Language Models through Activation Steering

### v1

@@ Appendix D Steering Layer Selection
We observe that for two instructions, “Constrained Response” and “Title,” no layer improves performance with steering. The “Constrained Response” instruction asks the model to respond using only “yes,” “no,” or “maybe,” while the “Title” instruction requires the answer to include a title wrapped in angular brackets (e.g., “<>absent<>< >”). In these cases, the steering vectors sometimes push the model too aggressively toward outputting instruction-related tokens, resulting in poor outputs (e.g., repeating characters like “<<<<much-less-thanabsentmuch-less-than<<<<< < < <”).
For language instructions, the selected layers are listed in Table 9. For length instructions, we intervene at layer 12 in Phi-3 and layer 16 in Gemma 2 9B. For word-specific instructions, we apply steering at layer 24 for both Phi-3 and Gemma 2 2B. These layers were chosen based on validation using a small set of manually created prompts.
Appendix E Response Quality Evaluation
E.1 Evaluation Procedure
To evaluate the quality of the model-generated responses, we first use the base queries outlined in Appendix B and prompt GPT-4o to produce 5 questions about the query to evaluate any model-generated answer to the query (prompt reported in Table 10). For each model-generated response in our evaluation, we then prompt GPT-4o to provide a binary “yes/no” answer to each question and “Not Applicable” if the question is not applicable to the query, along with a reason (prompt in Table 11). An aggregate quality score is computed as the average % of questions that is scored as “yes”. We run the experiment three times for each model and report the quality score and standard error to account for any non-determinism from the GPT-4o-based evaluator.
@@ E.1 Evaluation Procedure
The following is a prompt that is used to evaluate the generations from a large language model. We do not know how to evaluate the quality of model answers for this prompt. Can you come up with 5 or less questions that can break down the quality to simpler evaluation tasks that we can then ask about the model answer? Each question should have a simple yes, no answer.
@@
E.2 Examples of Quality Score Changes
@@
E.3 Additional Response Quality Results
@@
We apply the response quality evaluation procedure described in Appendix F to outputs generated by steering the model for length and word-specific instructions, following the same approach used for format constraints in § 3. As in the previous section, we analyze the outputs produced by Phi-3 Mini, reporting the score changes due to steering both with and without explicit text instructions, as well as the quality score differences when text instructions are added without steering.
@@
Length Instructions. Figure 9a presents the results for length constraints. We analyze a subset of 100 outputs from the generation process used for Figure 5b in § 4 (which evaluated 200 examples). The model is steered with a weight of $c=20$ on a set of base prompts, applying different length constraints requesting the model to answer using at most $n$ sentences ($n\in{1,\dots,5}$). The two leftmost columns in Figure 9a show the changes in the quality score when steering the model on inputs with and without length instructions. These values are averaged over the 5 constraints, as the steering procedure applied is the same way for all 5 constraints (adding the conciseness vector with weight $c=20$). A negative delta is expected, as steering the model to produce shorter responses generally reduces the comprehensiveness of the answer. This pattern is confirmed by the rightmost columns in Figure 9a, which represent the difference in quality score when length instructions are explicitly provided in the input text. The different colors correspond to the number of sentences ($n$), and as expected, shorter length constraints lead to larger drops in the quality score. Although the effect of steering is not directly comparable to any specific length constraint (since steering applies an additional reduction in length without a one-to-one mapping between steering weight and sentence count), we observe that the changes in quality scores due to steering remain within a reasonable range when compared to those induced by text instructions.
@@
Word-specific Instructions. In Figures 9b and 9c, we report the changes in quality scores for instructions that request the inclusion or exclusion of a specific word in the answer. In this case, introducing text instructions on top of the base query does not result in a noticeable drop in quality. Similarly, steering the model toward word inclusion/exclusion without explicit text instructions does not significantly impact quality (as seen from the near-zero deltas in the leftmost columns). However, when explicit instructions are provided in the prompt, we observe a slight decrease in the quality score, around 0.05. To further investigate this, we examine the empirical distributions of the score deltas for word inclusion instructions (Figure 10). The distributions appear centered around 0, with a few outliers showing deviations larger than 0.4. We perform a paired two-sided t-test to assess whether the means of these three distributions are significantly different. The p-values for the three tests (one for each pair) are all greater than 0.05 (0.12, 0.06, and 0.30). While the lack of significance does not prove that the distributions are identical, we interpret this as evidence of the minimal impact our steering procedure has on the model’s ability to address the base queries.
@@
Appendix F Implementation and Experimental Details
[…]
Steering Weight Selection. We base the selection of the steering weight on the computation in Eq. 2. For format instructions, the weight from Eq. 2 is used directly. For length and format instructions, the value computed in Eq. 2 is averaged across different inputs and is used as a reference to determine a range of suitable steering weights. For length instructions, we experiment with multiple weight values, highlighting the continuous nature of the constraint. For word inclusion instructions, we notice that in some cases (particularly with Gemma 2B), the weight is too low to impact the model’s output. This could be because the steering vector is averaged across inputs regardless of whether the model satisfies the constraint. This issue tends to be more pronounced for word-specific constraints (than, e.g., for format instructions) since the strength of the instruction signal can vary greatly depending on whether the word is related to the base query (e.g., including a word related to the query is easier than including an unrelated one). To address this, we perform a small additional grid search on a set of held-out examples, refining the weight based on the initial value from Eq. 2. For word exclusion instructions, we perform the grid search around the negative of the value computed for word inclusion.
Tools and Libraries. For length-related experiments, word counts are measured as the number of space-separated sequences of characters and sentence counts are determined using NLTK (Bird & Loper, 2004). To assess whether the score improvement from steering is significantly different from standard inference, we carry out McNemar’s test (McNemar, 1947). In particular, we use the exact version of the test, which uses the binomial distribution and is more conservative. The error bars reported in Figures 4 and 9 represent the standard error of the quality score computed over three different runs of GPT-4o. The smoothed empirical distributions shown in Figures 5a, 5c, 8c, and 12 are obtained using kernel density estimation (Parzen, 1962). Our experiments were carried out using PyTorch (Paszke et al., 2019) and the TransformersLens library (Nanda & Bloom, 2022). We performed our data analysis using NumPy (Harris et al., 2020) and Pandas (Wes McKinney, 2010). Our figures were made using Plotly (Plotly Technologies Inc., 2015). The paper’s bibliography was curated using Ryanize-bib (Zouhar, 2023).
Figure 12: Output Length Distribution with Verbosity Steering. The distribution of output lengths when steering Phi-3 using a vector for verbosity, with varying steering weights.
Appendix G Additional Results
@@
Length Instructions: Steering for Longer Outputs. In Figure 12, we show the distribution of output lengths when the model is steered using a verbosity vector. This vector is computed from inputs containing instructions to generate verbose responses (e.g., “Provide a long answer”). Similar to our procedure for the conciseness vector in § 4, we apply different steering weights to modulate verbosity. Using the same 50 base prompts as in Figure 5a, the results show that higher steering weights result in longer outputs, demonstrating that activation steering can effectively control output length by adjusting a single scalar value.

### v2

@@ Appendix E Steering Layer and Weight Selection
Layer Selection Results. In Table 8, we provide the layer indices used in our steering procedure for each instruction in the format subset (excluding language) across all models and evaluation settings (with and without explicit text instructions). We observe that for “Constrained Response” and “Title,” virtually no layer improves performance with steering. The “Constrained Response” instruction asks the model to respond using only “yes,” “no,” or “maybe,” while the “Title” instruction requires the answer to include a title wrapped in angular brackets (e.g., “<>absent<>< >”). In these cases, the steering vectors sometimes push the model too aggressively toward outputting instruction-related tokens, resulting in poor outputs (e.g., repeating characters like “<<<<much-less-thanabsentmuch-less-than<<<<< < < <”). For language instructions, the selected layers are listed in Table 9. For length instructions, we intervene at layer 12 in Phi-3 and layer 16 in Gemma 2 9B. For word-specific instructions, we use Phi-3’s layer 24 and 28, and Gemma 2 2B IT’s layers 24 and 22 for for inclusion and exclusion, respectively. The validation results used for layer and weight selection selection are reported in Appendix K).
Candidate Weights. We base the selection of the steering weight on the computation in Eq. 2. For format instructions, the weight from Eq. 2 is used directly (i.e., for each layer, there procedure above include a single layer-weight combination). For length and format instructions, the value computed in Eq. 2 is averaged across different inputs and is used as a reference to determine a range of suitable steering weights. For length instructions, we experiment with multiple weight values, highlighting the continuous nature of the constraint. For word inclusion instructions, we notice that in some cases (particularly with Gemma 2B), the weight is too low to impact the model’s output. This could be because the steering vector is averaged across inputs regardless of whether the model satisfies the constraint. This issue tends to be more pronounced for word-specific constraints (than, e.g., for format instructions) since the strength of the instruction signal can vary greatly depending on whether the word is related to the base query (e.g., including a word related to the query is easier than including an unrelated one). To address this, in the selection procedure described above, we consider a small set of weights based on the value $\bar{c}$ (i.e., the value $c$ computed as in Eq. 2 averaged across different inputs). For Phi-3 $\bar{c}\approx 52$ at layer 28 and $\bar{c}\approx 42$ at layer 26, and we use the set $\{40,60,80,100\}$ as grid-search value across the layers $\{24,26,28\}$. For Gemma 2 2B IT, $\bar{c}\approx 55$ at layer 22 and $\bar{c}\approx 74$ at layer 24, and we use the set $\{60,80,100,120\}$ as grid-search value across the layers $\{22,24\}$.
@@
The following is a prompt that is used to evaluate the generations from a large language model. We do not know how to evaluate the quality of model answers for this prompt. Can you come up with 5 of less questions that can break down the quality to simpler evaluation tasks that we can then ask about the model answer? Each question should have a simple yes, no answer.
@@
Appendix F Response Quality Evaluation
F.1 Evaluation Procedure
To evaluate the quality of the model generated responses, we first use the base queries outlined in Appendix C and prompt GPT-4o to produce 5 questions about the query to evaluate any model generated answer to the query (prompt reported in Table 10). For each model generated response in our evaluation, we then prompt GPT-4o to provide a binary “yes/no” answer to each question and “Not Applicable” if the question is not applicable to the query, along with a reason (prompt in Table 11). An aggregate quality score is computed as the average % of questions that is scored as “yes”. We run the experiment three times for each model and report the quality score and standard error to account for any non-determinism from the GPT-4o based evaluator.
@@ F.1 Evaluation Procedure
F.2 Examples of Quality Score Changes
@@
Figure 11: Quality score deltas for format instructions across all models and settings. (a) Phi-3, (b) Gemma 2 2B IT, (c) Mistral 7B IT, (d) Gemma 2 9B IT. Bars represent the average quality score delta under three conditions: steering without explicit instructions (purple), steering with explicit instructions (orange), and no steering but comparing outputs with and without explicit instructions (light blue).
F.3 Additional Results
@@
We apply the response quality evaluation procedure to outputs generated by steering the model for length and word-specific instructions, following the same approach used for format constraints in § 3. As in the previous section, we analyze the outputs produced by Phi-3, reporting the score changes due to steering both with and without explicit text instructions, as well as the quality score differences when text instructions are added without steering.
@@
Length Instructions. Figure 9a presents the results for length constraints. We analyze a subset of 100 outputs from the generation process used for Figure 5b in § 4 (which evaluated 200 examples). The model is steered with a weight of $c=20$ on a set of base prompts, applying different length constraints requesting the model to answer using at most $n$ sentences ($n\in{1,\dots,5}$). The two leftmost columns in Figure 9a show the changes in quality score when steering the model on inputs with and without length instructions. These values are averaged over the 5 constraints, as the steering procedure applied is the same way for all 5 constraints (adding the conciseness vector with weight $c=20$). A negative delta is expected, as steering the model to produce shorter responses generally reduces the comprehensiveness of the answer. This pattern is confirmed by the rightmost columns in Figure 9a, which represent the difference in quality score when length instructions are explicitly provided in the input text. The different colors correspond to the number of sentences ($n$), and as expected, shorter length constraints lead to larger drops in quality score. Although the effect of steering is not directly comparable to any specific length constraint (since steering applies an additional reduction in length without a one-to-one mapping between steering weight and sentence count), we observe that the changes in quality scores due to steering remain within a reasonable range when compared to those induced by text instructions.
Word-specific Instructions. In Figures 9b and 9c, we report the changes in quality scores for instructions that request the inclusion or exclusion of a specific word in the answer.
@@
The decrease in quality is minimal across all scenarios, the largest drop observed when steering for word inclusion without explicit instructions (around -0.03). We further investigate this by examining the empirical distributions of the score deltas for word inclusion instructions (Figure 10). The distributions appear centered around 0, with a few outliers showing deviations larger than 0.4. We perform a paired two-sided t-test to assess whether the means of these three distributions are significantly different. The p-values for the three tests (one for each pair) are all greater than 0.05 (0.25, 0.98, and 0.26). While the lack of significance does not prove that the distributions are identical, we interpret this as evidence of the minimal impact our steering procedure has on the model’s ability to address the base queries.
Format: Breakdown by Instruction. Figure 11 presents the quality score deltas for format instructions, broken down by individual instructions across all four models (Phi-3, Gemma 2 2B IT, Mistral 7B IT, and Gemma 2 9B IT). The three settings compared are steering without explicit instructions, steering with explicit instructions, and no steering but comparing outputs with and without explicit instructions.
F.4 Quality Score/Accuracy Trade-off
We investigate how the steering weight $c$ mediates the trade-off between instruction-following accuracy and output quality degradation for word exclusion constraints. Figure 12 shows the quality score and accuracy obtained by steering Phi-3 with varying negative values of $c$ (as we subtract the word-inclusion vector, as explained in § 5 and Appendix J). In both settings (with and without explicit text instructions), larger absolute steering weights lead to higher instruction-following accuracy. This improvement comes at the cost of a gradual decrease in quality score, indicating a trade-off between strict adherence to constraints and maintaining output comprehensiveness. The smooth transitions in both metrics suggest that tuning $c$ offers fine-grained control over this balance, highlighting the flexibility of activation steering.
Figure 12: Trade-off Between Quality Score and Accuracy on Word Exclusion. (a) Results for steering without explicit text instructions, showing increasing accuracy at the cost of a gradual decrease in quality. (b) Results for steering with explicit instructions, demonstrating a similar trend with higher baseline accuracy.
Appendix G Implementation and Experimental Details
[…]
Tools and Libraries. For length-related experiments, word counts are measured as the number of space-separated sequences of characters and sentence counts are determined using NLTK (Bird & Loper, 2004). To assess whether the score improvement from steering is significantly different from standard inference, we carry out McNemar’s test (McNemar, 1947). In particular, we use the exact version of the test, which uses the binomial distribution and is more conservative. The error bars reported in Figures 4 and 9 represent the standard error of the quality score computed over three different runs of GPT-4o. The smoothed empirical distributions shown in Figures Figure 5a, Figure 5c, Figure 8c, Figure 17a, and Figure 17c are obtained using kernel density estimation (Parzen, 1962). Our experiments were carried out using PyTorch (Paszke et al., 2019) and the TransformersLens library (Nanda & Bloom, 2022). We performed our data analysis using NumPy (Harris et al., 2020) and Pandas (Wes McKinney, 2010). Our figures were made using Plotly (Plotly Technologies Inc., 2015). The paper’s bibliography was curated using Ryanize-bib (Zouhar, 2023).
@@ Appendix G Implementation and Experimental Details
Appendix H Additional Results: Representations
@@
Figure 14: t-SNE visualization of per-example instruction vectors. (a) Format instruction vectors at layer 20 of Phi-3 for format instructions. Distinct clusters emerge for certain instructions (e.g., “No Comma,” “Lowercase,” “JSON Format”), while others (e.g., “End Checker,” “Constrained Response”) show less clear separation. (b) Vectors derived from length instructions at layer 12 of Phi-3. Vectors corresponding to concise outputs cluster on one end, while those for larger sentence counts ($n>2$) cluster on the opposite end.
Geometry of Steering Vectors. We perform a t-SNE dimensionality reduction on per-example instruction vectors, calculated as the difference in activations between inputs with and without explicit instructions, at layer 20 of Phi-3. The visualization for format instructions shows varying levels of cluster separation (Figure 14a). Specifically, instructions like “No Comma,” “Lowercase,” “JSON Format,” “Title,” and “Quotation” form distinct and well-separated clusters, indicating the model’s ability to encode these constraints clearly. Instructions such as “Capitalize” and “Capital Word Frequency” cluster closely together, reflecting their similar semantic nature. In contrast, vectors for instructions like “End Checker” and “Constrained Response” are more dispersed, with no clear clustering, suggesting these instructions are less distinctly represented in the model’s activation space.
Figure 14b presents a t-SNE dimensionality reduction of vectors computed for length instructions at layer 12 of Phi-3. Specifically, we include vectors derived from sentence-specific instructions (e.g., “answer using $n$ sentences,” $n\in{1,\dots,5}$) as well as more general “conciseness” and “verbosity” instructions, as described in § 4. A clear linear trend emerges, with vectors for longer output instructions clustering at one end of the plot and those for concise outputs clustering at the opposite end.
Figure 15: Accuracy Breakdown by Format Instruction Across Models and Settings. (a) Phi-3, (b) Gemma 2 2B IT, (c) Mistral 7B IT, and (d) Gemma 2 9B IT. Bars represent accuracy for format instructions under four settings: no explicit instruction (blue), steering without explicit instruction (red), explicit instruction without steering (green), and explicit instruction with steering (purple).
Figure 16: Results from the layer selection procedure. (a, b) Examples from Phi-3 in the no-instruction setting: (a) for “Multiple Sections,” the perplexity check prevents selecting a layer with degraded output quality; (b) for “Lowercase,” accuracy improves without quality loss. (c) For Gemma 2 2B IT with explicit “JSON Format” instructions, later layers lead to over-steering, with a degradation in both accuracy and quality.
Appendix I Additional Results: Format Instructions
Validation Results. Figure 16 provides examples of the results from the layer selection procedure. As detailed in Appendix E, we evaluate instruction-following accuracy and the fraction of low-perplexity outputs after steering at different layers. Panels (a) and (b) show examples obtained with Phi-3 in the no-instruction setting. In the first case (“Multiple Sections” instruction), the perplexity check prevents the selection of a layer that would otherwise lead to a decrease in output quality, as indicated by a sharp rise in the fraction of low-perplexity outputs. Panel (c), which corresponds to the “JSON Format” instruction applied to Gemma 2 2B IT with explicit input instructions, illustrates a scenario where accuracy and output quality are correlated: steering at later layers degrades both quality and accuracy.
Accuracy Breakdown by Instruction. Figure 15 provides a breakdown of instruction-following accuracy across specific format instructions for all models and evaluation settings: no explicit instruction, steering without explicit instruction, explicit instruction without steering, and explicit instruction with steering. Across all models, steering improves accuracy for many instructions in the absence of explicit text instructions (blue vs. red bars). Notably, casing-related instructions (“Lowercase” and “English Capital”) benefit strongly from steering, with Phi-3 and Gemma 2 2B IT achieving higher accuracy with steering alone than with explicit instructions (red vs. green bars). Certain instructions like “Title” and “Constrained Response” show near-perfect adherence with explicit instructions alone, leaving little room for improvement through steering.
Figure 17: Additional Results for Length Instructions. (a) The distribution of output lengths when steering Phi-3 using a vector for verbosity, with varying steering weights. (b) Steering significantly improves accuracy for sentence-specific length constraints ($n\in{1,\dots,5}$) in 4 of 5 cases. (c) For $n=4$, steering shifts the output length distribution closer to the target.
Appendix J Additional Results: Length Instructions
@@ Appendix I Additional Results: Format Instructions
Steering for Longer Outputs. In Figure 17a, we show the distribution of output lengths when the model is steered using a verbosity vector. This vector is computed from inputs containing instructions to generate verbose responses (e.g., “Provide a long answer”). Similar to our procedure for the conciseness vector in § 4, we apply different steering weights to modulate verbosity. Using the same 50 base prompts as in Figure 5a, the results show that higher steering weights result in longer outputs, demonstrating that activation steering can effectively control output length by adjusting a single scalar value.
Steering for Exact Length. We experiment with sentence-specific length instructions to evaluate the effectiveness of steering in improving adherence to explicit length constraints. Specifically, we compute steering vectors for instructions such as “Answer using $n$ sentences” ($n\in{1,\dots,5}$) and apply them to steer the model toward satisfying the corresponding constraints. In these experiments, we use Phi-3 and the same combination of steering layer and weight as in § 4 (layer 12, $c=20$). The results in Figure 17b show that steering significantly improves the model’s adherence to the specified lengths when explicit instructions are present in the input. Notably, steering produces statistically significant improvements in 4 out of the 5 cases. Figure 17c illustrates the effect of steering on the distribution of output lengths for a specific constraint ($n=4$ sentences). Steering shifts the distribution closer to the target length, with a sharper peak at the desired value, while still maintaining some natural variability.
Figure 18: Steering Layer and Weight Selection for Word-specific Instructions. Each panel shows instruction-following accuracy (x-axis) vs. fraction of low-perplexity outputs (y-axis). Blue dots represent validation runs; red markers indicate the selected configuration with maximum accuracy and zero low-perplexity outputs. Results are shown for Phi-3 and Gemma 2 2B IT.
Table 13: Examples of Tokens Promoted by Word-exclusion Vectors. Projecting word-exclusion vectors onto the vocabulary results in cases where the excluded word is promoted. This suggests that word-exclusion vectors may not effectively steer the model toward avoiding these words. The vectors are computed at layer 30 of Phi-3.
Word Top Tokens
books _books, _Books, books, _book, engl
congress _Congress, _parlament, _Legisl
Europe _European, _europé, Europe, _Europe
urbanization _urban, urban, _Urban, _urb, _rural
chromosomes chrom, _chrom, engl, _genom, _Zob
water vapor _water, _Water, water, _agua
fish _fish, _Fish, fish, _aqu, _marine
adult _youth, _Youth, _Child, _Children
hypothesis _instead, engl, _invece, _Instead
Figure 19: Validation Accuracy vs. Perplexity Trade-off for Word Exclusion. Subtracting word-inclusion vectors (green) outperforms adding word-exclusion vectors (red), showing the limitations of the latter.
Appendix K Additional Results: Word-specific Instructions
Validation Results. As described in Appendix E, to determine the optimal steering layer and weight parameter for each model-setting combination, we validate our approach by jointly considering instruction-following accuracy and the fraction of low-perplexity outputs. The x-axis in each panel of Figure 18 represents the instruction-following accuracy, while the y-axis shows the fraction of low-perplexity outputs, which we aim to minimize. For each model and setting, we select the layer and weight combination that achieves the highest accuracy, provided the fraction of low-perplexity outputs is zero. Red markers indicate the chosen configurations.
Steering using Word-exclusion Vectors. Following our approach for word inclusion, we compute instruction vectors for word exclusion. However, projecting these exclusion vectors onto the model’s vocabulary via the unembedding matrix often results in high logit values for tokens corresponding to the word that should be excluded. Examples of this phenomenon are shown in Appendix J. While this effect does not occur in all cases (e.g., for words like “adult” and “hypothesis”), the presence of a positive signal for the excluded word in the logit space suggests that these vectors may counteract the intended steering effect. To confirm this, we compare the performance of steering using word-exclusion vectors versus subtracting word-inclusion vectors for the same words. As shown in Figure 19, subtracting word-inclusion vectors consistently yields higher accuracy and achieves a better accuracy-to-low-perplexity trade-off than adding word-exclusion vectors. These results indicate that subtracting inclusion vectors is a more effective and reliable method for steering the model to avoid specific words.